A 2.0-kg mass is projected from the edge of the top of a 20-m tall building with a velocity of 24 m/s at some unknown angle above the horizontal. Disregard air resistance and assume the ground is level. What is the speed of the mass just before it strikes the ground?

a 42 m/s

b 31m/s

c 51 m/s

d 27 m/s

e none of the above
Answer #1

v^2 -u^2 = 2as

v2 = u2 +2as = (24)2 + (2*9.8*20) =968

v = (968)1/2 = 31 m/s

Option (b) is correct